OpenStudy (anonymous):

What is the slope of the line? y = (1/3)x + 2 A. -2 B. -1/3 C. 1/3 D. 2

OpenStudy (shaik0124):

y=mx+c compare this with y=1/3x+2 then we will get m=1/3

OpenStudy (shaik0124):

where m is slope

OpenStudy (shaik0124):

did u understood